Suppose you have 6 types of cereal. If you eat a bowl of cereal every morning for breakfast, how many ways can you eat breakfast in the next 4 days?
Answer by stanbon(75887)

Suppose you have 6 types of cereal. If you eat a bowl of cereal every morning for breakfast, how many ways can you eat breakfast in the next 4 days?
----
Assuming you have only one bowl of each cereal you get:
1st day: 6 choices
2nd day: 5 choices
3rd day: 4 choices
4th day: 3 choices
-----
Ans: 6*5*4*3 = 30*12 = 360 choices
-------------
OR
Ans: 6P4 = 360
========================
Cheers,
Stan H.
======================